OpenStudy (anonymous):

Calculate the molarity of a solution made by dissolving 4.75 g of HCl in enough water to form exactly 750 mL of solution.

OpenStudy (anonymous):

\[\frac{ weight of solute \times 1000 }{ molecular weight of solute \times volume (ml) }\] \[ molarity=\frac{4.75 \times 1000 }{ 36.5 \times 750}\]